OpenStudy (anonymous):

6. What is the value of log16 4? (1 point) (the 16 is small and at the bottom of log) 4 1/2 1/4 2 7. Solve the exponential equation. 125^9x – 2 = 150 (1 point) –1.8847 –0.1069 0.3375 1.0378 8. Solve log(3x + 2) = 3. (1 point) 1000/3 998/3 998 1/3

OpenStudy (anonymous):

6. For subscripts you can use "_". Log_16 (4) = x means that: 16^x = 4 (4^2)^x = 4 4^(2x) = 4 2x = 1 x=1/2 7. Please use brackets I'm not sure what the question is exactly 8. What's the base for the log here?

OpenStudy (anonymous):

7.) 125^(9x-2)=150 8.) log(3x+2)=3

OpenStudy (anonymous):

7. 125^(9x-2)=150 means that: log_125 (150) = 9x-2 log_125 (150) +2 = 9x (log_125 (150) +2)/9 = x 8. You haven't changed anything from the original equation

OpenStudy (anonymous):

that's cus that is thee equation.

OpenStudy (anonymous):

your answer is one of the choses this is where im lost.